# Examining the Expanding Role of Synthetic Data Throughout the AI Development Pipeline

### v1

@@ 6. Conclusion
In this paper, we illustrate the expanding use of synthetic data in AI development. Through interviews with AI practitioners and RAI experts, we highlight the motivations and promises of synthetic data, its various use cases across the AI development pipeline, current practices and challenges around its generation and validation, its limitations, and ethical considerations for its use. We discuss implications of our findings for the AI supply chain. Finally, we propose several considerations for improving synthetic data practices. We note that while synthetic data is valued precisely for its efficiency, practices like rigorous validation, documentation, or participatory methods all require human intervention, increasing costs and extending timelines. To arrive at adoptable practices, we must reconcile competing demands between the need for rigor and the desire to maintain efficiency in line with organizational priorities and constraints.
@@ Acknowledgements.
First and foremost, we are grateful to our study participants, without whom this research would not be possible. Our sincere thanks to Hanna Wallach and Ece Kamar for making connections to the AI practitioner community and helping with recruitment. We are deeply thankful to Sachita Nishal, Anubhav Jangra, and all members of the Microsoft Research FATE group for their support and feedback throughout the study. Finally, we thank Inha Cha, Jordan Taylor, and Sarah Fox for feedback on early drafts of the work.
@@ A.2.1. Introduction
Scenario 1: Scenario 2: Scenario 3:
Fine-tuning data Evaluation Simulated data
Table 2. Vignettes of Synthetic Data Use

### v2

@@ 6. Conclusion
In this paper, we illustrate the expanding role of synthetic data in AI development. Through interviews with AI practitioners and RAI experts, we highlight the motivations and promises of synthetic data, its various use cases across the AI development pipeline, current practices and challenges around its generation and validation, its limitations, and ethical considerations for its use. We discuss the implications of our findings for the AI supply chain. Finally, we propose several considerations for improving synthetic data practices. We note that while synthetic data is valued precisely for its efficiency, practices like rigorous validation, documentation, or participatory methods all require human intervention, increasing costs and extending timelines. To arrive at adoptable practices, we must reconcile competing demands between the need for rigor and the desire to maintain efficiency in line with organizational priorities and constraints.
7. End Matter
Ethical considerations. We took several measures to protect our research participants. Our study design was reviewed and approved by our organization’s Institutional Review Board. Participation was voluntary. Informed consent was obtained electronically and we compensated participants with a gift card of \$75 USD. Participants were informed of the purpose of the study and were told that they could refuse to answer any questions and ask for the recording to be paused at any time. All participant data, including interview transcripts and recordings, was stored in a secure location and was not accessible by anyone outside the research team. The research team checked all included quotes to ensure that no identifying information about participants or their organizations was revealed.
Positionality. All authors are, or were at the time of conducting the research, employed at a large technology company. This positioning afforded us access to ongoing conversations within the industry around synthetic data practices, including emerging use cases. At the same time, we recognize that our proximity to industry priorities may shape how we frame risks and recommendations. Throughout this work, we have aimed to remain accountable to perspectives that may not align with institutional interests by foregrounding scholarship from critical data studies and the political economy of the AI supply chain. We offer this reflection to acknowledge how our embeddedness within industry shapes the kinds of questions we ask and the interpretations we bring to our analysis, and to invite ongoing dialogue with communities who approach this topic from different positions or commitments.
Potential adverse impacts. First, our findings focus exclusively on practitioners’ and RAI experts’ perspectives. Our findings may inadvertently privilege their viewpoints and concerns over those of other crucial stakeholders. This could lead to recommendations that undervalue the needs of affected communities, policymakers, and civil society organizations. Second, our research sheds light on motivations, practices, and challenges for the use of synthetic data in the AI development pipeline. We also expose limitations and ethical considerations. Emphasizing these drawbacks may be interpreted as evidence that the use of synthetic data should be more restricted. However, we caution against this interpretation of our results. In setting policy guidance for synthetic data, policymakers should weigh the benefits as well. Perhaps most notably, synthetic data enables the evaluation of models and systems at a scale that would not be possible through other means; without it, many evaluations simply may not happen. To address this concern, we have aimed throughout this paper to provide a balanced perspective, highlighting both the benefits and drawbacks of synthetic data. We additionally emphasize that norms and practices around the use of synthetic data are rapidly evolving. While we hope that our research inspires future work developing best practices for the responsible use of synthetic data, these practices will need to be constantly evaluated and updated to ensure that they remain practical and relevant.
@@ Acknowledgements.
First and foremost, we are grateful to our study participants, without whom this research would not be possible. This work was enriched by many useful conversations with Anubhav Jangra, Nari Johnson, Daniela Massiceti, Cecily Morrison, Sachita Nishal, and the members of the Microsoft Research FATE group, Microsoft’s Sociotechnical Alignment Center, and the Tech Solidarity Lab at CMU. Our sincere thanks to Hanna Wallach and Ece Kamar for making connections to the AI practitioner community and helping with recruitment. We are also grateful to Agathe Balayn, Inha Cha, Sarah Fox, and Jordan Taylor for providing feedback on early drafts of this paper.
@@ A.2. Interview Protocol and Vignettes for Phase 2
Scenario 1: Scenario 2: Scenario 3:
Fine-tuning data Evaluation Simulated data
Table 2. Vignettes of Synthetic Data Use